# Safe Leaf Manipulation for Accurate Shape and Pose Estimation of Occluded Fruits

### v1

@@ II-A Autonomous Agriculture
There is a growing research trend towards enhancing autonomy in precision agriculture. Approaches include designing specialized hardware [34], developing autonomous systems [29], and advancing perception algorithms [30]. Among these, the most relevant literature for this work pertains to active perception algorithms that reduce occlusions by moving the camera to gather informative observations based on viewpoint planning methods [41]. Zaenker et al[38, 39, 40] propose viewpoint planning methods for maximizing the coverage of the region of interests, e.g., fruit areas. Burusa et al[4, 3] utilize attention-driven and semantic-aware view planning for reconstruction and detection of relevant plant parts. However, these will not handle heavy occlusions, as the robot might need to physically move aside dense foliage to observe the fruits.
@@ II-B Interactive Perception
Although some authors have attempted to simulate plant deformation [7], deformation prediction is challenging because the structure of the plant is unknown due to occlusion, and material properties are not practical to measure in the field. In the context of manipulation action planning on plants, existing methods avoid constructing a precise deformation model by either employing self-supervised learning [42] or pushing on branches [10]. Zhang and Gupta [42] suggest a data-driven push planning method for plant space revealing. However, their method requires extensive time for collecting training data for a particular scenario (over 30 hours for vines and over 18 hours for Dracaena), which makes it difficult to generalize to unseen scenarios and plants. Additionally, their approach focuses solely on space revealing rather than addressing the task of fruit monitoring. Gursoy et al[10] propose a branch push planning method for occlusion handling in fruit detection. However, their method considers just ellipses for fruit shape estimation. Additionally, in most cases, fruit occlusion is primarily caused by leaves, which cannot be easily resolved by simply moving branches. For our setting, we use an energy-based deformation graph model [33] that is constructed from the semantic segmentation of the observed point cloud.
@@ III-A Scene-Consistent Shape Completion (SSC)
The key to our scene-consistent shape completion (SSC) is incorporating peduncle information. We assume that the peduncle is attached to the branch that may also be occluded by leaves. To address this issue, we first segment the scene from an RGB-D input and perform branch completion before integrating the data into our fruit shape completion method.
@@ III-A1 Semantic Segmentation
We segment the RGB image using a pre-trained open-vocabulary segmentation model, GroundedSAM [28]. As this model may suffer from a noisy background, we exclude RGB pixels with a large depth value ($>1.5$ m in our setup) before feeding them into the model. We use the text prompt [branch, leaf, sweet pepper] to segment the corresponding parts of the plant into instances. The segmented pixels are de-projected into 3D space using the camera intrinsic and depth image to obtain the semantic point clouds, denoted as $P_{\text{branch}}$, $P_{\text{leaf}}$, and $P_{\text{fruit}}$.
@@ III-A3 Fruit Shape Completion
$\mathcal{L}_{\text{neg}}=\frac{1}{\left|P_{\text{neg}}\right|}\sum_{{p}^{w}\in P_{\text{neg}}}\max(0,-D_{\theta}(T_{w\rightarrow o}{p}^{w},\mathrm{z}))$ (2)
@@
$\mathcal{L}_{\text{ped}}=\frac{1}{\left|P_{\text{ped}}\right|}\sum_{\left({p}^{o},{p}^{w}\right)\in\left(P_{\text{ped}},P_{\text{cor}}\right)}\left\|T_{w\rightarrow o}^{-1}{p}^{o}-{p}^{w}\right\|^{2}$ (3)
@@
$\mathcal{L}=w_{\text{sur}}\mathcal{L}_{\text{sur}}+w_{\text{neg}}\mathcal{L}_{\text{neg}}+w_{\text{ped}}\mathcal{L}_{\text{ped}}+w_{\text{reg}}\mathcal{L}_{\text{reg}}$ (4)
@@ III-B Perception-Based Deformation Model
Given the scene completed from perception data, we instantiate a leaf deformation model to predict its deformed states in response to a given robot action. An action $a=(p^{g},R^{g},\delta p^{g})$ is defined by a grasp pose $(p^{g},R^{g})$ and by a displacement vector $\delta p^{g}$. $p^{g}\in\mathbb{R}^{3}$ is the grasped point and $R^{g}\in SO(3)$ defines gripper orientation. With a pull direction and a certain move distance, the displacement $\delta p^{g}\in\mathbb{R}^{3}$ of the grasp point $p^{g}$ can be easily calculated.
@@
$p^{\prime}_{i}-p^{\prime}_{j}=R^{\prime}_{i}(p_{i}-p_{j}),\ \forall(i,j)\in\mathcal{E}$ (5)
@@
$E(P,P^{\prime})=\sum_{(i,j)\in\mathcal{E}}w_{ij}\cdot\|p^{\prime}_{i}-p^{\prime}_{j}-R^{\prime}_{i}(p_{i}-p_{j})\|^{2}$ (6)
@@
$\hat{P}^{\prime}$ $=\arg\min_{P^{\prime}}E(P,P^{\prime})$ (7)
@@
$p_{i}^{\prime}$ $=p_{i},\quad\quad\quad\forall p_{i}\in P_{fix}$
@@ III-C Leaf Manipulation Action Planning (LMAP)
The goal of our leaf manipulation action planning (LMAP) is to reduce the risk of damaging the plant by energy prediction $\hat{E}$ and maximize predicted fruit visibility based on deformation states $\hat{P}^{\prime}$, completed fruit mesh $M_{fruit}$ extracted from estimated $D_{\theta}$, and the robot mesh.
@@ III-C1 Action Generation and Collision Checking
Our LMAP method first samples action candidates on the edges of the leaf to avoid twisting the leaf. First, we project the $P_{leaf}$ from 3D to 2D using PCA. Then, we fit a convex hull on the 2D points where we can meaningfully define the leaf’s boundary. We sample 15 grasp points on the edges of the convex hull with a small margin of 1 cm towards the interior to enhance robustness. We uniformly sample 14 pull directions in the robot coordinate system, including 6 axial and 8 diagonal directions. We additionally consider action directions that form an angle of less than $120^{\circ}$ between the line connecting the grasped points and the direction of the branch to avoid pulling the leaf away from it. We sample move distances starting at 2 cm, increasing in 2 cm increments up to 8 cm, or until the predicted fruit visibility reaches 95% of its maximum (discussed in the next subsection). In this way, we obtain a set of action proposals denoted as $\mathcal{A}=\{a_{k}(p^{g}_{k},R^{g}_{k},\delta p^{g}_{k})\}$.
@@ III-C2 Visibility Prediction
We use the ray-casting of OctoMap (resolution 3 mm) from the camera viewpoint to count the number of fruit surfaces hit. The initial hits $H_{init}$ is computed by inserting $P_{fruit}$, $P_{branch}$, $P_{leaf}$ into OctoMap and counting the hits on $P_{fruit}$. The predicted hits $H_{k}$ of action $a_{k}$ is computed by inserting $M_{fruit}$, deformed states $\hat{P}^{\prime}_{k}$, robot mesh $M_{k}$ and counting on $M_{fruit}$. Hence, the predicted visibility of $a_{k}$ is computed by $N^{vis}_{k}=|H_{k}\setminus H_{init}|$, indicating the increased number of visible surfaces. The maximum visibility is computed by inserting only $M_{fruit}$.
@@ III-C3 Action Selection
$a_{k}^{*}=\arg\max_{a_{k}\in\mathcal{A}}(N^{vis}_{k}-\lambda\cdot\hat{E}_{k})$ (8)
@@ IV Experimental Results
IV-A Hardware Setup and Artificial Plant Scenarios
Our hardware setup uses a UR5 robot equipped with a Robotiq 2F-85 gripper and an calibrated eye-to-hand L515 RealSense camera. We tested 12 artificial plant scenarios, including two types of artificial leaves (small/large with thin/thick branches), three artificial peppers with varying shapes (circular and striped) and colors (red and yellow), and two different occlusion conditions. Representative scenarios are shown in Fig. 3(a) and (b). To have a quantitative evaluation, we collect the ground truth shape and pose of the artificial fruits and obtain 3D fruit models by the point-cloud scanning [25] with the mesh reconstruction [12]. For each scenario, the fruit pose is annotated through human labeling.
@@ IV-A Hardware Setup and Artificial Plant Scenarios
TABLE II: Evaluation of varied action planning methods on fake leaf and pepper testbed. The Random method is executed 5 times for each scenario. Visibility is normalized to the maximum, and mean and std. dev. are reported. Initial visibility is 0.44 $\pm$ 0.23. For unsafe trials, the visibility fraction is set to the initial value. Our method consistently plans safe manipulations while revealing the entire fruit.

### v2

@@ II-A Autonomous Agriculture
There is a growing research trend towards enhancing autonomy in precision agriculture. Approaches include designing specialized hardware [34], developing autonomous systems [29], and advancing perception algorithms [30]. Among these, the most relevant literature for this work pertains to active perception algorithms that reduce occlusions by moving the camera to gather informative observations based on viewpoint planning methods [43]. Zaenker et al[40, 41, 42] propose viewpoint planning methods for maximizing the coverage of the region of interests, e.g., fruit areas. Burusa et al[4, 3] utilize attention-driven and semantic-aware view planning for reconstruction and detection of relevant plant parts. However, these will not handle heavy occlusions, as there are no viewpoints that provide sufficient visibility, and the robot has to physically move the leaf aside.
@@ II-B Interactive Perception
The model-based approach requires a simulator of the plant; however, existing work is limited to predicting behavior with a perfect skeleton [7] or simulating contact forces [37, 38]. In the context of manipulation action planning on plants, existing methods avoid constructing a precise deformation model by either employing self-supervised learning [44] or pushing on branches [10]. Zhang and Gupta [44] suggest a data-driven push planning method for plant space revealing. However, their method requires extensive time for collecting training data for a particular scenario (over 30 hours for vines and over 18 hours for Dracaena), which makes it difficult to generalize to unseen scenarios and plants. Additionally, their approach focuses solely on space revealing rather than addressing the task of fruit monitoring. Gursoy et al[10] propose a branch push planning method for occlusion handling in fruit detection. However, their method considers just ellipses for fruit shape estimation. Additionally, in most cases, fruit occlusion is primarily caused by leaves, which cannot be easily resolved by simply moving branches. For our setting, we use an energy-based deformation graph model [33] that is constructed from the semantic segmentation of the observed point cloud.
@@ III-A Scene-Consistent Shape Completion (SSC)
The key to our scene-consistent shape completion (SSC) is incorporating free-space and peduncle information. We assume that the peduncle is attached to the branch that may also be occluded by leaves. To address this issue, we first segment the scene from an RGB-D input and perform branch completion before integrating the data into our fruit shape completion method.
@@ III-A1 Semantic Segmentation
We segment the RGB image using a pre-trained open-vocabulary segmentation model, GroundedSAM [28]. We exclude RGB pixels with a large depth value ($>1.5$ m in our setup) before feeding them into the model. We use the text prompt [branch, leaf, sweet pepper] to segment the corresponding parts of the plant into instances. The segmented pixels are de-projected into 3D space using the camera intrinsic and depth image to obtain the semantic point clouds, denoted as $P_{\text{branch}}$, $P_{\text{leaf}}$, and $P_{\text{fruit}}$.
@@ III-A3 Fruit Shape Completion
$\mathcal{L}_{\text{neg}}=\frac{1}{\left|P_{\text{neg}}\right|}\sum_{{p}^{w}\in P_{\text{neg}}}\max(0,-D_{\theta}(T_{w\rightarrow o}{p}^{w},\mathrm{z})).$ (2)
@@
$\mathcal{L}_{\text{ped}}=\frac{1}{\left|P_{\text{ped}}\right|}\sum_{\left({p}^{o},{p}^{w}\right)\in\left(P_{\text{ped}},P_{\text{cor}}\right)}\left\|T_{w\rightarrow o}^{-1}{p}^{o}-{p}^{w}\right\|^{2}.$ (3)
@@
$\mathcal{L}=w_{\text{sur}}\mathcal{L}_{\text{sur}}+w_{\text{neg}}\mathcal{L}_{\text{neg}}+w_{\text{ped}}\mathcal{L}_{\text{ped}}+w_{\text{reg}}\mathcal{L}_{\text{reg}}.$ (4)
@@ III-B Perception-Based Deformation Model
Given the completed scene, we instantiate a leaf deformation model to predict its deformed states in response to a given robot action. An action $a=(p^{g},R^{g},\delta p^{g})$ consists of a grasp pose $(p^{g},R^{g})$ and a displacement $\delta p^{g}$. We use grasping over pushing for more stable contact points [44], improving deformation precision. $p^{g}\in\mathbb{R}^{3}$ is the grasped point and $R^{g}\in SO(3)$ defines gripper orientation. $\delta p^{g}\in\mathbb{R}^{3}$ can be calculated using a pull direction and a move distance.
@@
$p^{\prime}_{i}-p^{\prime}_{j}=R^{\prime}_{i}(p_{i}-p_{j}),\ \forall(i,j)\in\mathcal{E}.$ (5)
@@
$E(P,P^{\prime})=\sum_{(i,j)\in\mathcal{E}}w_{ij}\cdot\|p^{\prime}_{i}-p^{\prime}_{j}-R^{\prime}_{i}(p_{i}-p_{j})\|^{2},$ (6)
@@
$\hat{P}^{\prime}$ $=\arg\min_{P^{\prime}}E(P,P^{\prime}),$ (7)
@@
$p_{i}^{\prime}$ $=p_{i},\quad\quad\quad\forall p_{i}\in P_{fix}.$
@@ III-C Leaf Manipulation Action Planning (LMAP)
The goal of our leaf manipulation action planning (LMAP) is to reduce the risk of damaging the plant by energy prediction $\hat{E}$ and maximize predicted fruit visibility based on deformation states $\hat{P}^{\prime}$, completed fruit mesh $M_{\text{fruit}}$ extracted from estimated $D_{\theta}$, and the robot mesh.
@@ III-C1 Action Generation and Collision Checking
Our LMAP method first samples action candidates on the edges of the leaf to avoid twisting the leaf. First, we project the $P_{\text{leaf}}$ from 3D to 2D using PCA. Then, we fit a convex hull on the 2D points where we can meaningfully define the leaf’s boundary. We sample 15 grasp points on the edges of the convex hull with a small margin of 1 cm towards the interior to enhance robustness. We uniformly sample 14 pull directions in the robot coordinate system, including 6 axial and 8 diagonal directions. We additionally consider action directions that form an angle of less than $120^{\circ}$ between the line connecting the grasped points and the direction of the branch to avoid pulling the leaf away from it. We sample move distances starting at 2 cm, increasing in 2 cm increments up to 8 cm, or until the predicted fruit visibility reaches 95% of its maximum (discussed in the next subsection). In this way, we obtain a set of action proposals denoted as $\mathcal{A}=\{a_{k}(p^{g}_{k},R^{g}_{k},\delta p^{g}_{k})\}$.
@@ III-C2 Visibility Prediction
We use the ray-casting of OctoMap (resolution 3 mm) from the camera viewpoint to count the number of fruit surfaces hit. The initial hits $H_{init}$ is computed by inserting $P_{\text{fruit}}$, $P_{\text{branch}}$, $P_{\text{leaf}}$ into OctoMap and counting the hits on $P_{\text{fruit}}$. The predicted hits $H_{k}$ of action $a_{k}$ is computed by inserting $M_{\text{fruit}}$, deformed states $\hat{P}^{\prime}_{k}$, robot mesh $M_{k}$ and counting on $M_{\text{fruit}}$. Hence, the predicted visibility of $a_{k}$ is computed by $N^{vis}_{k}=|H_{k}\setminus H_{init}|$, indicating the increased number of visible surfaces. The maximum visibility is computed by inserting only $M_{\text{fruit}}$.
@@ III-C3 Action Selection
$a_{k}^{*}=\arg\max_{a_{k}\in\mathcal{A}}(N^{vis}_{k}-\lambda\cdot\hat{E}_{k}),$ (8)
@@ IV Experimental Results
IV-A Hardware Setup and Testing Plant Scenarios
Our hardware setup uses a UR5 robot equipped with a Robotiq 2F-85 gripper and an calibrated eye-to-hand L515 RealSense camera. We tested 12 artificial and 16 real plant scenarios. The artificial scenarios included two types of artificial leaves (small/large with thin/thick branches), three artificial peppers with varying shapes (circular, striped) and colors (red, yellow), and two occlusion conditions as shown in Fig. 3 (a) and (b). The real scenarios involved leaves and peppers with natural variations as shown in Fig. 3 (c). For quantitative evaluation, we collect ground truth shapes and poses of artificial fruits using 3D point-cloud scanning [25] with mesh reconstruction [12]. The fruit pose for each scenario is annotated through human labeling.
@@ IV-A Hardware Setup and Testing Plant Scenarios
TABLE II: Evaluation of varied action planning methods on artificial scenarios. The Random method is executed 5 times for each scenario. Visibility is normalized to the maximum, and mean and std. dev. are reported. Initial visibility is 0.44 $\pm$ 0.23. For unsafe trials, the visibility fraction is set to the initial value. Our method consistently plans safe manipulations that fully reveal the fruit.